How many pounds of seed worth $0.60 a pound must be mixed with 300 pounds of seed worth $0.35 a pound to produce a mixture worth $0.50 per pound?


Thanks so much!!!

Answer by ankor@dixie-net.com(22740)

How many pounds of seed worth $0.60 a pound must be mixed with 300 pounds of seed worth $0.35 a pound to produce a mixture worth $0.50 per pound?
:
Let x = amt of .60/lb seed
:
.60x + .35(300) = .50(x + 300)
:
.60x + 105 = .50x + 150
:
.60x - .50x = 150 - 105
:
.1x = 45
:
x = 45/.1
:
x = 450 lb of .60/lb seed
:
:
Check solution:
.6(450) + .35(300) = .50(450 + 300)
270 + 105 = .5(750)
375 = 375